1. The Hsu Manufacturing Company has two service departments: Maintenance and Accounting. The Maintenance Department's costs...

1.

The Hsu Manufacturing Company has two service departments: Maintenance and Accounting. The Maintenance Department's costs of $792,000 are allocated on the basis of machine hours. The Accounting Department's costs of $171,600 are allocated on the basis of the number of employees within a specific department. The direct departmental costs for A and B are $250,000 and $450,000, respectively.

Maint Acctg A B
Machine hours 590 100 3,600 360
Number of employees 2 2 8 4


What is the Maintenance Department's cost allocated to Department A using the direct method?

2 and 3:

The Hsu Manufacturing Company has two service departments: Maintenance and Accounting. The Maintenance Department's costs of $489,600 are allocated on the basis of machine hours. The Accounting Department's costs of $122,400 are allocated on the basis of the number of employees within a specific department. The direct departmental costs for A and B are $210,000 and $410,000, respectively.

Maint Acctg A B
Machine hours 895 80 2,400 320
Number of employees 2 2 8 4


2. What is the Accounting Department's cost allocated to Department B using the direct method?

3. What is the Maintenance Department's cost allocated to Department B using the step method and assuming the Maintenance Department's costs are allocated first?
Answer #1

ANSWER NO 1

Maintenance Department Cost = $792000

cost allocated on the basis of machine hour

Department 'A' Allocation

=792000/(3600+360)*3600 = $720000

ANSWER NO 2

ACCOUNTING DEPARTMENT COST =$122400

Cost allocated on the basis of Number of employees

=122400/(8+4)*4 = $40800

ANSWER 3 step method

Maintenance department cost=$489600

it will allocated between accounting,A & B department

=489600/(80+2400+320)*320 = $55954